OpenStudy (anonymous):

The product of 2 consecutive positive integers is 5 more than their sum. Find the integers

OpenStudy (anonymous):

3 and 4. Sum = 7. Product = 12.

OpenStudy (anonymous):

Let the 2 integers be x and y xy = 5 + x + y since they are consecutive, let x be the first number and y be the second number. we can rewrite y as x+1(Consecutive integers) So our new equation is x(x + 1) = 5 + x + x + 1 x^2 + x = 6 + 2x x^2 - x - 6 = 0 (x + 2)(x - 3) = 0 So x = 3 and -2 SInce we are interested in the positive integer we go with, x = 3 so my y = 4 the two integers = 3 and 4